a boat traveled 88 miles downstream and back. the trip downstream took 4 hours. the trip back took 44 hours. what is the speed of the boat in still water? what is the speed of the current?
Answer by lwsshak3(11628)

a boat traveled 88 miles downstream and back. the trip downstream took 4 hours. the trip back took 44 hours. what is the speed of the boat in still water? what is the speed of the current?
**
let x=speed of boat in still water
let c=speed of current
x+c=net speed of boat downstream
x+c=net speed of boat upstream
travel time=distance/speed
..
88/(x+c)=4
88/(x-c)=44
..
4x+4c=88
44x-44c=88
..
multiply first equation by 11
44x+44c=968
44x-44c=88
add
88x=1056
x=12
4c=88-4x=88-48=40
c=10
ans:
speed of boat in still water=12 mph
speed of current=10 mph